A and B start at 7 a.m. from two towns 37.5 miles apart. A's rate of walking being twice B's. If they walk 5 hours before they meet. find their rates of walking.
Answer by mananth(16946)

Let B's rate be x
A's rate = 2x
They are waling towards each other. So their combined rate = 2x+x=3x
distance = 37.5 miles
time = 5
speed = distance / time
3x= 37.5/5
multiply equation by 5
15x= 37.5
/15
x= 37.5/15
x= 2.5 mph
B's walking rate = 2.5 mph
A's walking rate = 2.5*2 = 5 mph
m.ananth@hotmail.ca